# Riddle Me This! Stealthy Membership Inference for Retrieval-Augmented Generation

### v1

@@ 7.3. Financial Cost Analysis
Yes/No Question Generation. For this stage, we use GPT-4o to generate yes/no questions. Based on our analysis, the input to GPT-4o for this task is $902\pm 108$ tokens on average, and the output is $513\pm 64$ tokens on average. Based on these numbers, the cost for this stage is \$0.01 per document.
Description Generation. Similarly, for generating the description of each document, we use GPT-4o. Based on our analysis, the average number of input tokens is $648\pm 108$, and the average number of output tokens is $21\pm 5$. The cost for generating ground-truth answers is \$0.003 per document.
Ground-Truth Answer Generation. To generate the ground-truth answers, we use GPT-4o-mini. The average number of input tokens for this task is $13,244\pm 3,317$, and the average number of output tokens is $48\pm 5$. The cost for generating ground-truth answers is \$0.004 per document.
Based on these estimates, the total cost for processing each document is \$0.017.
@@ 8. Conclusion
Future Directions. In this work, we proposed a new black-box MIA against RAG systems, focusing on both attack success and detectability. While our attack consistently demonstrates high AUC scores across all settings and high TPR@low FPR in most cases, there are instances where its TPR is lower than one of the baselines. This indicates room for further improvement. Exploring more complex query forms, such as multiple-choice questions, could enhance the performance and serves as a promising direction for future work. Additionally, we evaluated our attack in a realistic setting where the RAG system rewrites the input query. Other variations of RAG systems, which involve different forms of input modification, remain unexplored. Extending evaluations to such settings would provide a broader understanding of the attack’s effectiveness and robustness.
@@ Appendix D RAG Without Query-Rewriting
Appendix E Prompts for Experimental Stages
In this section, we document the exact prompts used at various stages of our experimental setup. The prompt used to deploy GPT-4o as a prompt injection detector, including detailed instructions and examples, is presented in Figure 11. The few-shot prompt used to generate 30 yes/no questions with GPT-4o is shown in Figure 12. Following question generation, the prompt for generating the general description of each target document with GPT-4o is provided in Figure 14. Additionally, the short prompt for rewriting the input query of the RAG system is illustrated in Figure 13. This prompt is a modified version of the best-performing prompt reported in (Kirchenbauer et al., 2023). Finally, the RAG system prompt and the prompt used to generate ground-truth answers are presented in Figures 15 and 16, respectively.
Appendix F Failed Cases Examples
@@
Appendix G ROC Curves
For completeness, we provide ROC curves across all attacks and datasets for all of our experiments. These ROC curves are presented in Figures 19, 20, 21, and 22.

### v2

@@ 7.3. Financial Cost Analysis
Yes/No Question Generation. For this stage, we use GPT-4o to generate yes/no questions. Based on our analysis, the input to GPT-4o for this task is $902\pm 108$ tokens on average, and the output is $513\pm 64$ tokens on average. Based on these numbers, the cost for this stage is \$0.01 per document, taking an average of 6.86 s to run.
Description Generation. Similarly, for generating the description of each document, we use GPT-4o. Based on our analysis, the average number of input tokens is $648\pm 108$, and the average number of output tokens is $21\pm 5$. The cost for generating ground-truth answers is \$0.003 per document, taking an average of 0.51 s to run.
Ground-Truth Answer Generation. To generate the ground-truth answers, we use GPT-4o-mini. The average number of input tokens for this task is $13,244\pm 3,317$, and the average number of output tokens is $48\pm 5$. The cost for generating ground-truth answers is \$0.004 per document, taking an average of 0.48 s to run.
Based on these estimates, the total cost for processing each document is \$0.017, taking an average of 7.85 s to run. Of course, not all constraints are financial; some may be computational. In such cases, an adversary might opt for non-LLM approaches, such as rule-based question templates or human-in-the-loop systems for question and answer generation, trading flexibility and scale for lower resource demands.
@@ 8. Conclusion
Future Directions. In this work, we proposed a new black-box MIA against RAG systems, focusing on both attack success and detectability. While our attack consistently demonstrates high AUC scores across all settings and high TPR@low FPR in most cases, there are instances where its TPR is lower than one of the baselines. This indicates room for further improvement. Additionally, we evaluated our attack in a realistic setting where the RAG system rewrites the input query. Other variations of RAG systems, which involve different forms of input modification, remain unexplored. Extending evaluations to such settings would provide a broader understanding of the attack’s effectiveness and robustness.
@@ Appendix A Details for Detection Setup
We also include the exact attack queries for our attack (IA) and three baseline attacks (RAG-MIA, S2-MIA, MBA for a fixed document in Table 4.
Attack Query
Target Document This paper serves as a companion or extension to the “Inside PageRank” paper by Bianchini et al. [19]. It is a comprehensive survey of all issues associated with PageRank, covering the basic PageRank model, available and recommended solution methods, storage issues, existence, uniqueness, and convergence properties, possible alterations to the basic model, suggested alternatives to the traditional solution methods, sensitivity and conditioning, and finally the updating problem. We introduce a few new results, provide an extensive reference list, and speculate about exciting areas of future research.
[…]
Interrogation Attack (Ours) I want to ask about Comprehensive survey of PageRank issues, models, solution methods, and future research areas. Does the paper act as a companion or extension to the “Inside PageRank” paper by Bianchini et al.? Please answer with “Yes,” “No,” or “I don’t know”
Table 4. Example exact queries generated by each of the baseline attacks (RAG-MIA, S2-MIA, MBA and our attack (IA) for the same target document (before query rewriting).
@@ Appendix D RAG Without Query-Rewriting
Appendix E Making Prior Works Stealthy
Existing membership inference attacks against RAG models were developed assuming a plain, minimal RAG setup, without accounting for the additional components often present in real-world systems, such as detection mechanisms. This makes retrofitting stealth into these attacks difficult, as meaningful changes can start to resemble entirely new attacks. In this appendix, we attempt to modify standard attacks for improved stealth, but find that they still remain easily detectable.
E.1. Reducing Prompt Length via Document Summaries
Prior work uses the target document with minimal transformation (Anderson et al., 2024; Li et al., 2024c). In contrast, our attack relies on a combination of document summaries and questions based on distinct pieces of information from the document (as described in Section 5.1). As a result, the reduced prompt size may make our queries less conspicuous. To test whether it this reduced query length that leads to low detection rates, we modify existing attacks to use a summarized version of the document instead of the full document. Detection rates remain high even with the paraphrased document, as shown in Table 6. This outcome is expected, since the detection methods focus on identifying malicious intent in the query’s structure, not on the specific document content used.
Attack GPT4 Lakera
RAG-MIA (Anderson et al., 2024) 1.000 0.000
S2-MIA (Li et al., 2024c) 0.656 1.000
MBA (Liu et al., 2024b) 1.000 0.944
Table 6. Evaluating the effectiveness of malicious-input detection against baseline attacks for membership inference.
E.2. Adapting Queries to Elicit Binary Responses
While RAG-MIA already constructs queries that correspond to a ‘yes’ or ‘no’ response, other baselines like S2-MIA rely on the RAG system ‘autocompleting’ some given text. We therefore experiment with a variant of the S2-MIA attack by adapting the prompt to explicitly ask for a ‘yes’ or ‘no’ response. Specifically, for a document $d=d_{1}||d_{2}$, where $d_{1}$ and $d_{2}$ are the two halves, we modify the query to require a yes/no answer, as illustrated in Figure 12.
[…]
Figure 12. Attack Prompt for the modified S2-MIA attack, for some document $d=d_{1}||d_{2}$.
We find that this modified variant is still easily detected, with a detection score of 0.998 using a GPT-4 based classifier. However, Lakera detection scores decrease slightly, dropping from 0.070 to 0.036 for this modified variant. This is notable, as the Lakera classifier appears less effective at detecting queries framed as yes/no questions, consistent with the lower detection scores observed for both RAG-MIA and the modified S2-MIA attack.
Appendix F Prompts for Experimental Stages
In this section, we document the exact prompts used at various stages of our experimental setup. The prompt used to deploy GPT-4o as a prompt injection detector, including detailed instructions and examples, is presented in Figure 13. The few-shot prompt used to generate 30 yes/no questions with GPT-4o is shown in Figure 14. Following question generation, the prompt for generating the general description of each target document with GPT-4o is provided in Figure 16. Additionally, the short prompt for rewriting the input query of the RAG system is illustrated in Figure 15. This prompt is a modified version of the best-performing prompt reported in (Kirchenbauer et al., 2023). Finally, the RAG system prompt and the prompt used to generate ground-truth answers are presented in Figures 17 and 18, respectively.
Appendix G Failed Cases Examples
@@
Appendix H ROC Curves
For completeness, we provide ROC curves across all attacks and datasets for all of our experiments. These ROC curves are presented in Figures 21, 22, 23, and 24.